Consider the following hypotheses tests involving the χ2-distribution.

(a) Determine the p-value for Ho: P(1) = P(2) = P(3) = P(4) = 0.25, with χ2 = 6.53. (Give your answer bounds exactly.)
< p <

(b) Determine the p-value for Ho: P(I) = 0.25, P(II) = 0.40, P(III) = 0.35, with χ2 = 7.21. (Give your answer bounds exactly.)
< p <
Answer #1

Solution :

(a)

df = 4 - 1 = 3

P-value = P(3 > 6.53) = 0.0885

0.05 < p < 0.10

(b)

df = 3 - 1 = 2

P-value = P(2 > 7.21) = 0.0272

0.025 < P-value < 0.05